# From Sim-to-Real: Toward General Event-based Low-light Frame Interpolation with Per-scene Optimization

### v1

@@ 4 Method
Due to the trailing artifacts and other degradations in events that are difficult to be modeled at the training stage, we will introduce a per-scene optimization method in Sec. 4.1, with its associated loss function presented in Sec. 4.2. Moreover, to obtain a better per-trained model, we also present improvements to the low-light event simulation process in Sec. 4.3. The overall pipeline is summarized in Fig. 3.
4.1 Per-scene Optimization for VFI
@@
Unlike previous Event-VFI methods [14, 34], which rely solely on two consecutive noisy RGB frames and associated event signals to predict intermediate RGB frames, our per-scene optimization strategy leverages the entire sequence. This approach, depicted in Fig. 3, allows for fine-tuning a pre-trained Event-VFI model using the entire video sequence. It establishes training pairs directly from the testing sequence, using the denoised RGB image $I_{t}$ at time $t$ as the ground truth to supervise the interpolation results from two other captured RGB frames ($I_{t-n}$ and $I_{t+n}$) at neighboring timestamps, along with event signals ($E_{t-n,t}$ and $E_{t,t+n}$) captured during this interval, where $n$ represents the temporal interval between captured RGB frames and is randomly sampled to enrich the data for per-scene optimization.
After per-scene fine-tuning, during the actual inference stage, the network utilizes the same sequence to generate a new frame $\hat{I}_{t+\Delta t}$, effectively increasing the frame rate.
4.2 Optimization Loss
In the per-scene optimization process, the interpolated frame at timestamp $t$, $\hat{I}_{t}$, can be modeled as:
$\hat{I}_{t}=\Phi(D(I_{t-n}),D(I_{t+n}),E_{t-n,t},E_{t,t+n}),$ (8)
where $\Phi(\cdot)$ is the event-based frame interpolation network, $D(\cdot)$ is the real-world denoising operator, and $\lambda$ is a weighting parameter which we set to 0.1. In our experiment, we choose state-of-the-art Event-VFI model TimeLens [30] as interpolation backbone $\Phi(\cdot)$ and Restormer [31] as denoiser $D(\cdot)$. Then the reconstruction loss can be defined as:
$\mathcal{L}_{r}=\sqrt{\|\hat{I}_{t}-D(I_{t})\|^{2}+\epsilon^{2}}+\lambda\sqrt{\|S(\hat{I}_{t})-S(D(I_{t}))\|^{2}+\epsilon^{2}},$ (9)
where $\lambda$ is set as 0.1, $S(\cdot)$ is the Sobel operator , $\sqrt{\|\hat{x}-x\|^{2}+\epsilon^{2}}$ is the Charbonnier penalty function, and $\epsilon$ is set to 0.001. Additionally, to obtain visually pleasing interpolation results, the perceptual loss [10] is also used:
$\mathcal{L}_{p}=\beta\sqrt{\|\text{VGG}(\hat{I}_{t})-\text{VGG}(D(I_{t}))\|^{2}+\epsilon^{2}},$ (10)
where $\text{VGG}(\cdot)$ is the pretrained 16 layer VGG network [23], $\beta$ is a weighting parameter which we set to 0.1.
4.3 Low Light Event Simulation
@@
Figure 5: Synthesis process for the EVS motion trail simulation dataset.
@@
Fig. 5 shows the proposed simulation pipeline. First, we create a high-frame linear RGB sequences using the GoPro dataset [32]. Initially, we employ RIFE [7] to generate high frame rate video. Subsequently, we compute a low light intensity image using $L=\Gamma^{-1}(I)$, where $\Gamma^{-1}(\cdot)$ represents the inverse of the image signal processing (ISP) pipeline. In applying $\Gamma^{-1}(\cdot)$, we follow [3] and incorporate inverse tone mapping, inverse gamma correction, inverse color correction matrix, and inverse white balance and brightening to transition the data to a linear brightness domain.
@@
In the second step, to create events from high-frame-rate linear RGB sequences, we use v2e simulator to produce a synthetic RGB-EVS dataset. Under low-light conditions, the degradation of event cameras can generally be attributed to motion trails and noise. The v2e [6] simulator comprehensively models noise factors through parameter adjustments, such as threshold noise, hot pixels, leak noise events, and refractory period violations. To more accurately simulate motion trailing associated with events, we employ a second-order low-pass filter to approximate the RC circuit dynamics described in Eqn. 3 of Sec. 3.1.
In the last step, we further improve the quality of generated event signals. Even we increase the input video frame rate using RIFE interpolation method, the resulting video is still a temporal discretization of the event signal. Event signals often occur randomly within brief intervals between frames. To mitigate this issue, we have refined the low-pass filter approach by incorporating a sparse event mask $R$ that resets certain event pixels $I_{t-\Delta t}$ to zero, allowing these pixels to respond exclusively to instantaneous light intensity changes in the current frame. This adjustment helps induce sparsity in the event tails, better approximating real-world scenarios. Consequently, we reformulate the low-pass filter equation in Eqn. 3 as follows:
$I^{\prime}_{t}=R\alpha I_{t}+(1-R\alpha)I_{t-\Delta t},$ (11)
@@
where $R$ is a random binary mask, $\alpha$ is the event trailing factor as shown in Eqn. 3. The event-RGB data pairs generated from this modified filter are utilized to pretrain the frame interpolation model. By employing the designed simulation pipeline, our method achieves superior results compared to the original TimeLens, as demonstrated in Tab. 1 in low light Event-VFI.
@@
Table 1: Comparison of different frame interpolation methods on EVFI-LL under various interpolation rates (4$\times$ and 8$\times$) and different bias voltages. The lower the bias, the lower the threshold, the more events recorded. The Sony IMX636 sensor in the Prophesee EVK4-HD event camera has positive and negative threshold voltages biased to 0 by default.
@@
+20 SuperSloMo [8] 30.201 0.8802 0.1371 27.665 0.8559 0.1828
RIFE [7] 31.143 0.8846 0.1420 28.523 0.8642 0.1854
TimeLens [30] 30.548 0.8673 0.1691 29.187 0.8559 0.1887
CBMNet [11] 30.114 0.8592 0.1999 28.656 0.8465 0.2246
@@
0 SuperSloMo [8] 31.105 0.8857 0.1264 28.724 0.8626 0.1632
RIFE [7] 32.242 0.8901 0.1301 29.462 0.8687 0.1706
TimeLens [30] 30.729 0.8647 0.1584 29.891 0.8542 0.1728
CBMNet [11] 30.323 0.8662 0.1878 29.320 0.8552 0.2100
@@
-20 SuperSloMo [8] 30.980 0.8880 0.1269 28.912 0.8683 0.1648
RIFE [7] 32.061 0.8920 0.1307 29.702 0.8744 0.1689
TimeLens [30] 31.417 0.8743 0.1572 30.541 0.8662 0.1706
CBMNet [11] 30.838 0.8607 0.1892 29.814 0.8509 0.2120
@@
5 Experiments
5.1 Experiment Details
We selected the state-of-the-art TimeLens [30] as the backbone for our Event-VFI model. Initially, a model tailored for low-light Event-VFI, referred to as Our Pretrained, is trained using simulated data (as detailed in Sec. 4.3). Before inference phase for each scene, per-scene optimization is first performed using the Adam optimizer [12] with a learning rate of $1\times 10^{-4}$. The optimized model is denoted as Our Per Scene Opt. It is important to note that to simplify the learning process, we pre-denoise the input RGB images using Restormer [31].
Our method is compared with last RGB-VFI techniques (SuperSloMo [8] and RIFE [7]) and state-of-the-art Event-VFI methods (CBMNet [11] and TimeLens [30]).
5.2 Low Light Event-VFI Benchmark: EVFI-LL
Previous real-world Event-VFI datasets, i.e., BS-ERGB [30] and ERF-X170FPS [11], were captured under normal lighting conditions and are not suitable for testing the trailing artifacts in low light events. Therefore, we have captured a new low-light Event-VFI benchmark, denoted as EVFI-LL.
Similar to ERF-X170FPS [11], our setup employs a beam splitter to align the fields of view of the event and RGB cameras, as depicted in Fig. 6. We use a Prophesee EVK4-HD ($1280\times 720$) event camera and a MER2-301-125U3C ($2048\times 1536$) RGB camera. Our system ensures that both cameras are geometrically calibrated and temporally synchronized. Calibration is achieved by estimating transformations using a blinking checkerboard pattern, and synchronization is managed with synchronization trigger controls.
Figure 6: EVS-RGB beam splitting imaging system. Spatial-temporal alignment is ensured through calibration and trigger control.
@@ 5.2 Low Light Event-VFI Benchmark: EVFI-LL
For testing purposes, we subsampled the test video by skipping every N-1 frames, which resulted in an effective testing frame rate reduced by a factor of N. Event data is not skipped during testing. This lower frame rate poses a significant challenge for frame interpolation tasks, using the skipped frames as ground truth for performance evaluation. We also curated more challenging scenes with complex and large motions within the EVFI-LL dataset, creating a challenging track denoted as EVFI-LL-C.
5.3 Results on Real-world Low-Light Event-VFI
We evaluated methods on EVFI-LL dataset at 4$\times$ and 8$\times$ interpolation rates across various event triggering thresholds. The quantitative results are presented in Tab. 1. For evaluation, we employed both numerical metrics, such as PSNR and SSIM, and the perceptual metric, LPIPS [33]. We also conducted evaluations on the challenging track, EVFI-LL-C, with quantitative comparisons detailed in Tab. 2. Visual comparisons are shown in Fig. 7.
In Tabs. 1 and 2, one interesting observation is that when the interpolated ratio is small (4×4$\times$4 ×), event-based methods (TimeLens and CBMNet) even perform worse than RGB-only methods (SuperSloMo and RIFE). This is because the degradation of the low-light event signal and the domain gap between training and testing significantly hurt the performance of event-based solution. Still, in more challenging 8×8$\times$8 × interpolation, the Event-VFI methods performs better than RGB-VFI methods, because the additional event stream does provide more information about intermediate motion. Despite this, as shown in Fig. 7, the previous methods fail to account for the trailing artifacts of low-light events, leading to visually unpleasing interpolated results with noticeable deformations around object edges.
@@ 5.3 Results on Real-world Low-Light Event-VFI
Table 2: Comparison of different frame interpolation methods in challenging track in EVFI-LL-C with 8$\times$ interpolation rates.
@@
SuperSloMo [8] 26.269 0.8039 0.2341
RIFE [7] 26.668 0.8122 0.2437
TimeLens [30] 30.945 0.8401 0.1818
CBMNet [11] 30.552 0.8431 0.2357
@@
Figure 7: Visual comparison of results from different methods under large motion.
Figure 8: Visual comparison of the low-light trailing simulation results of the event camera with v2e[6].
Figure 9: Visual comparison of ablation study results.
6 Discussion
@@
Low light simulation pipeline. In Tab. 3, it is evident that Our Pretrained (TimeLens) achieves an improvement of 0.5 to 1 dB over the original TimeLens under low-light conditions, demonstrating the effectiveness of our simulation pipeline. Fig. 8 offers visual comparisons of our simulation with typical v2e simulation[6]. The visualizations distinctly show the sparse and asymmetric event patterns, which are characteristic of low-light conditions and the fundamental operating principles of event cameras. Unlike merely adjusting the parameters of the v2e simulator, our method synthesizes event representations that convincingly mimic real sensor data, including accurate depictions of the elongated event streaks caused by latency.
@@
Per-scene optimization. We propose a per-scene optimization strategy to address the challenges of Event-VFI under low-light conditions. By creating supervision pairs directly from the test sequence, the Event-VFI network can be optimized to incrementally adapt to the current scene’s specific event degradation. Our experimental results, detailed in Tab. 3, show that our per-scene optimization strategy, Our Per Scene Opt. (Timelens), significantly enhances interpolation quality under real-world low-light conditions compared to the unoptimized model, Our Pretrained (Timelens). As illustrated in Fig. 1, the optimized outcomes reveal sharply recovered motion details, and observe no ghosting or trailing artifacts.
@@
Running time of per-scene optimization. We conducted a comparative analysis of the running times for the per-scene optimization stage and the inference stage on an NVIDIA RTX 4090. Given that the optimization process encompasses data loading durations, we calculated the total runtime of the whole program. For the configuration where eight frames are interpolated between two images of size $736\times 576$, TimeLens exhibited an approximate average runtime of 0.16 seconds, whereas the per-scene optimization process required about average 0.22 seconds. Notably, even though our method incorporates a training component, the time taken for per-scene optimization is significantly less than the time required for training from scratch, which takes approximately three days to train models like Our Pretrained (TimeLens). The duration of the per-scene optimization is roughly on the same order of magnitude as the inference time.
Generalize to other backbones. To demonstrate the generalizability of per-scene optimization across different backbones, we also applied such strategy to CBMNet. As shown in Tab. 3, Our Per Scene Opt. (CBMNet) exhibited a performance enhancement of $\sim$1.8 dB in low-light conditions. The visual comparison in Fig. 9 evident that the Our Per Scene Opt. (CBMNet) effectively reduces the impact of event trailing artifacts, yielding visually pleasing results. This improvement underscores the potential of per-scene optimization to significantly enhance the generalizability of other Event-VFI algorithms. Additionally, since the original CBMNet model was not trained considering low-light trailing effects, the performance of Our Per Scene Opt. (CBMNet) were somewhat inferior to those achieved with Our Per Scene Opt. (TimeLens), highlighting the importance of a well-pretrained model.
Table 3: Ablation Study on ×$\times 8$ Frame Interpolation on EVFI-LL
@@
7 Conclusion

### v2

@@ 4. Method
Due to the trailing artifacts and other degradations in events that are difficult to model during training, we will introduce a per-scene optimization method in Sec. 4.1, with its associated loss function presented in Sec. 4.2. Moreover, to obtain a better per-trained model, we present improvements to the low-light event simulation process in Sec. 4.3. The overall pipeline is summarized in Fig. 3.
4.1. Per-scene Optimization for VFI
@@
Unlike previous Event-VFI methods (Liang et al., 2023; Zhou et al., 2023), which rely solely on two consecutive noisy RGB frames and associated event signals to predict intermediate RGB frames, our per-scene optimization strategy leverages the entire sequence. This approach, depicted in Fig. 3, allows for fine-tuning a pre-trained Event-VFI model using the entire video sequence. It establishes training pairs directly from the testing sequence, using the denoised RGB image $I_{t}$ at time $t$ as the ground truth to supervise the interpolation results from two other captured RGB frames ($I_{t-n}$ and $I_{t+n}$) at neighboring timestamps, along with event signals ($E_{t-n,t}$ and $E_{t,t+n}$) captured during this interval. Here, $n$ represents the temporal interval between captured RGB frames and is randomly sampled to enrich the data for per-scene optimization. Specifically, in our experiments, the values of $n$ are sampled from 1 to 7 with corresponding probabilities $[0.632,0.232,0.086,0.032,0.012,0.004,0.002]$.
After per-scene fine-tuning, during the actual inference stage, the network utilizes the same sequence to generate a novel frame $\hat{I}_{t+\Delta t}$, effectively increasing the frame rate.
4.2. Optimization Loss
In the per-scene optimization process, the interpolated frame at timestamp $t$, $\hat{I}_{t}$ can be modeled as:
(8) $\hat{I}_{t}=\Phi(D(I_{t-n}),D(I_{t+n}),E_{t-n,t},E_{t,t+n}),$
where $\Phi(\cdot)$ is the event-based frame interpolation network, $D(\cdot)$ is the real-world denoising operator. In our experiment, we choose state-of-the-art Event-VFI model TimeLens (Tulyakov et al., 2021) as interpolation backbone $\Phi(\cdot)$ and Restormer (Zamir et al., 2022) as denoiser $D(\cdot)$. Then the reconstruction loss can be defined as:
(9) $\mathcal{L}_{r}=\sqrt{\|\hat{I}_{t}-D(I_{t})\|^{2}+\epsilon^{2}}+\lambda\sqrt{\|S(\hat{I}_{t})-S(D(I_{t}))\|^{2}+\epsilon^{2}},$
where $\lambda$ is a weighting parameter set to 0.1, $S(\cdot)$ is the Sobel operator, $\sqrt{\|\hat{x}-x\|^{2}+\epsilon^{2}}$ is the Charbonnier penalty function, and $\epsilon$ is set to 0.001. Additionally, to obtain visually pleasing interpolation results, the perceptual loss (Johnson et al., 2016) is also used:
(10) $\mathcal{L}_{p}=\beta\sqrt{\|\text{VGG}(\hat{I}_{t})-\text{VGG}(D(I_{t}))\|^{2}+\epsilon^{2}},$
where $\text{VGG}(\cdot)$ is the pretrained 16 layer VGG network (Russakovsky et al., 2015), $\beta$ is a weighting parameter which we set to 0.1.
4.3. Low Light Event Simulation
@@
Figure 5. Synthesis process for the EVS Motion Trail Simulation Dataset.
@@
Fig. 5 shows the proposed simulation pipeline. First, we create high-frame RGB sequences using the GoPro dataset (Zhang et al., 2018a). Then, we employ RIFE (Huang et al., 2022) for 8$\times$ frame interpolation to generate high frame rate video. Subsequently, we compute a low light intensity image using $L=\Gamma^{-1}(I)$, where $\Gamma^{-1}(\cdot)$ represents the inverse of the image signal processing (ISP) pipeline. Following (Brooks et al., 2019), $\Gamma^{-1}(\cdot)$ contains inverse tone mapping, inverse gamma correction, inverse color correction, and inverse white balance and brightening to transition the data to a linear brightness domain.
@@
Then, we generate events from high-frame-rate linear RGB sequences using the v2e simulator to produce a synthetic EVS-RGB dataset. Under low-light conditions, the performance degradation of event cameras is generally due to motion trails and noise. For the event noise, the v2e simulator (Hu et al., 2021) comprehensively models noise factors through parameter adjustments, such as threshold noise, hot pixels, leak noise events, and the refractory period. Before creating simulated data, we calibrated v2e’s parameters by comparing simulated data with real-world data. To accurately simulate the motion trailing associated with events, we use a low-pass filter with exponential decay to approximate the RC circuit dynamics described in Eqn. 3 in Sec. 3.1. Moreover, although we increase the input video frame rate using the RIFE interpolation method, the resulting video still remains a temporal discretization $\Delta t$ of the event signal. To model temporal signal decay within $\Delta t$ and modeling sparse trailing artifacts, we incorporate a sparse event mask $R$ in the low-pass filter that resets certain event pixels $I_{t-\Delta t}$ to zero. As a result, we reformulate the low-pass filter equation in Eqn. 3 into an Intensity-Dependent Stochastic Low-Pass Filter (IDS-LPF), as shown below:
(11) $I^{\prime}_{t}=R(\alpha,r)I_{t}+(1-R(\alpha,r))I_{t-\Delta t},$
@@
where $R(\alpha,r)=\alpha\cdot\mathbb{I}(r>P)+1\cdot\mathbb{I}(r\leq P)$. $r$ is a randomly generated number between 0 and 1, $\mathbb{I}(\cdot)$ is the indicator function that returns 1 if the condition is true and 0 otherwise. In our simulation, $P$ is set to 0.5. Consequently, the indicator function $\mathbb{I}(r>P)$ becomes to 0 when the randomly generated value $r$ is greater than $P$. The visual effects of applying the IDS-LPF have been shown in Fig. 6 to better illustrate the process of tailing simulation. The event-RGB data pairs generated from this modified filter are utilized to pretrain the frame interpolation model. By employing this simulation pipeline, our method achieves superior results compared to the original TimeLens, as demonstrated in Tab. 1 for low-light Event-VFI.
Figure 6. Visual comparison of the low-light trailing simulation results of the event camera with v2e.
@@
Table 1. Comparison of different frame interpolation methods on EVFI-LL under various interpolation rates (4$\times$ and 8$\times$) and different bias voltages (Sony IMX636 event sensor, positive and negative threshold voltages with a default bias of 0). The lower the bias, the lower the threshold, the more events recorded.
@@
+20 SuperSloMo (Jiang et al., 2018) 30.201 0.8802 0.1371 27.665 0.8559 0.1828
RIFE (Huang et al., 2022) 31.143 0.8846 0.1420 28.523 0.8642 0.1854
TimeLens (Tulyakov et al., 2021) 30.548 0.8673 0.1691 29.187 0.8559 0.1887
CBMNet (Kim et al., 2023) 30.114 0.8592 0.1999 28.656 0.8465 0.2246
@@
0 SuperSloMo (Jiang et al., 2018) 31.105 0.8857 0.1264 28.724 0.8626 0.1632
RIFE (Huang et al., 2022) 32.242 0.8901 0.1301 29.462 0.8687 0.1706
TimeLens (Tulyakov et al., 2021) 30.729 0.8647 0.1584 29.891 0.8542 0.1728
CBMNet (Kim et al., 2023) 30.323 0.8662 0.1878 29.320 0.8552 0.2100
@@
-20 SuperSloMo (Jiang et al., 2018) 30.980 0.8880 0.1269 28.912 0.8683 0.1648
RIFE (Huang et al., 2022) 32.061 0.8920 0.1307 29.702 0.8744 0.1689
TimeLens (Tulyakov et al., 2021) 31.417 0.8743 0.1572 30.541 0.8662 0.1706
CBMNet (Kim et al., 2023) 30.838 0.8607 0.1892 29.814 0.8509 0.2120
@@
5. Experiments
5.1. Experiment Details
We selected the state-of-the-art TimeLens (Tulyakov et al., 2021) as the backbone for our Event-VFI model. Initially, a model tailored for low-light Event-VFI, referred to as Our Pretrained, is trained using simulated data (as detailed in Sec. 4.3). Before inference phase for each scene, per-scene optimization is first performed using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $1\times 10^{-4}$. The optimized model is denoted as Our Per Scene Opt. It is important to note that to simplify the learning process, we pre-denoise the input RGB images using Restormer (Zamir et al., 2022).
Our method is compared with last RGB-VFI techniques (SuperSloMo (Jiang et al., 2018) and RIFE (Huang et al., 2022)) and state-of-the-art Event-VFI methods (CBMNet (Kim et al., 2023) and TimeLens (Tulyakov et al., 2021)).
5.2. Low Light Event-VFI Benchmark: EVFI-LL
Previous real-world Event-VFI datasets, i.e., BS-ERGB (Tulyakov et al., 2021) and ERF-X170FPS (Kim et al., 2023), were captured under normal lighting conditions and are not suitable for testing the trailing artifacts in low light events. Therefore, we have captured a new low-light Event-VFI benchmark, denoted as EVFI-LL.
Similar to ERF-X170FPS (Kim et al., 2023), our setup employs a beam splitter to align the fields of view of the event and RGB cameras, as depicted in Fig. 7. We use a Prophesee EVK4-HD ($1280\times 720$) event camera and a MER2-301-125U3C ($2048\times 1536$) RGB camera. Our system ensures that both cameras are geometrically calibrated and temporally synchronized. Calibration is achieved by estimating transformations using a blinking checkerboard pattern, and synchronization is managed with synchronization trigger controls.
@@ 5.2. Low Light Event-VFI Benchmark: EVFI-LL
For testing purposes, we subsampled the test video by skipping every N-1 frames, which resulted in an effective testing frame rate reduced by a factor of N. Event data is not skipped during testing. This lower frame rate poses a significant challenge for frame interpolation tasks, using the skipped frames (with denoising applied) as ground truth for performance evaluation. We also curated more challenging scenes with complex and large motions within the EVFI-LL dataset, creating a challenging track denoted as EVFI-LL-C.
Figure 7. EVS-RGB Beam Splitting Imaging System. Spatial-temporal alignment is ensured through calibration and trigger control.
Figure 8. Visual comparison of results from different methods under large motion.
Figure 9. Visual comparison of ablation study results.
5.3. Evaluation on Real-world Low-Light Event-VFI
We evaluated methods on EVFI-LL dataset at 4$\times$ and 8$\times$ interpolation rates across various event triggering thresholds. The quantitative results are presented in Tab. 1. For evaluation, we employed both numerical metrics, such as PSNR and SSIM, and the perceptual metric, LPIPS (Zhang et al., 2018b). We also conducted evaluations on the challenging track, EVFI-LL-C, with quantitative comparisons detailed in Tab. 2. Visual comparisons are shown in Fig. 8.
In Tabs. 1 and 2, one interesting observation is that when the interpolated ratio is small (4×4$\times$4 ×), event-based methods (i.e., TimeLens and CBMNet) even perform worse than RGB-only methods (i.e., SuperSloMo and RIFE). This is because the degradation of the low-light event signal and the domain gap between simulated event and real-world event significantly hurt the performance of event-based solution. Still, in more challenging 8×8$\times$8 × interpolation, the Event-VFI methods performs better than RGB-VFI methods, because the additional event stream does provide more information about intermediate motion. Despite this, as shown in Fig. 8, the previous methods fail to account for the trailing artifacts of low-light events, leading to visually unpleasing interpolated results with noticeable deformations around object edges.
@@ 5.3. Evaluation on Real-world Low-Light Event-VFI
Table 2. Comparison of Different Frame Interpolation Methods in challenging track in EVFI-LL-C with 8$\times$ Interpolation Rates.
@@
SuperSloMo (Jiang et al., 2018) 26.269 0.8039 0.2341
RIFE (Huang et al., 2022) 26.668 0.8122 0.2437
TimeLens (Tulyakov et al., 2021) 30.945 0.8401 0.1818
CBMNet (Kim et al., 2023) 30.552 0.8431 0.2357
@@
5.4. Discussion
Low light simulation pipeline.
@@ 5.4. Discussion
In Tab. 3, it is evident that Our Pretrained (TimeLens) achieves an improvement of 0.5 to 1 dB over the original TimeLens under low-light conditions, demonstrating the effectiveness of our simulation pipeline. Fig. 6 offers visual comparisons of our simulation with typical v2e simulation. The visualizations distinctly show the sparse and asymmetric event patterns, which are characteristic of low-light conditions and the fundamental operating principles of event cameras. Unlike merely adjusting the parameters of the v2e simulator, our method synthesizes event representations that convincingly mimic real sensor data, including accurate depictions of the elongated event streaks caused by latency.
Per-scene optimization.
@@
We propose a per-scene optimization strategy to address the challenges of Event-VFI under low-light conditions. By creating supervision pairs directly from the test sequence, the Event-VFI network can be optimized to incrementally adapt to the current scene’s specific event degradation. Our experimental results, detailed in Tab. 3, show that our per-scene optimization strategy, Our Per Scene Opt. (TimeLens), significantly enhances interpolation quality under real-world low-light conditions compared to the unoptimized model, Our Pretrained (TimeLens). As illustrated in Fig. 1, the optimized outcomes reveal sharply recovered motion details, and observe no ghosting or trailing artifacts.
Running time of per-scene optimization.
@@
We conducted a comparative analysis of the running times for the per-scene optimization stage and the inference stage on an NVIDIA RTX 4090. Given that the optimization process encompasses data loading durations, we calculated the total runtime of the whole program. For the configuration where eight frames are interpolated between two images of size $736\times 576$, TimeLens exhibited an approximate average runtime of 0.16 seconds, whereas the per-scene optimization process required about average 0.22 seconds. Notably, even though our method incorporates a training component, the time taken for per-scene optimization is significantly less than the time required for training from scratch, which takes approximately three days to train models like Our Pretrained (TimeLens). The duration of the per-scene optimization is roughly on the same order of magnitude as the inference time.
Table 3. Ablation Study on 8$\times$ Frame Interpolation on EVFI-LL
@@
Generalize to other backbones.
To demonstrate the generalizability of per-scene optimization across different backbones, we also applied such strategy to CBMNet. As shown in Tab. 3, Our Per Scene Opt. (CBMNet) exhibited a performance enhancement of $\sim$1.8 dB in low-light conditions. The visual comparison in Fig. 9 evident that the Our Per Scene Opt. (CBMNet) effectively reduces the impact of event trailing artifacts, yielding visually pleasing results. This improvement underscores the potential of per-scene optimization to significantly enhance the generalizability of other Event-VFI algorithms. Additionally, since the original CBMNet model was not trained considering low-light trailing effects, the performance of Our Per Scene Opt. (CBMNet) were somewhat inferior to those achieved with Our Per Scene Opt. (TimeLens), highlighting the importance of a well-pretrained model.
Non-reference evaluation.
Due to low-light conditions and camera limitations, capturing noise-free ground truth images is difficult. However, despite the denoising process, the positions of contents in the ”ground truth” images (used in Tab. 1) remain accurate. To further minimize the denoising algorithm’s impact on evaluation, we used non-reference image quality metrics like MUSIQ (Ke et al., 2021) in Tab. 4. Our interpolation results, with fewer artifacts, achieved superior MUSIQ evaluations.
Table 4. Non-reference Quality Evaluation on the EVFI-LL-C Dataset
Method RIFE TimeLens CBMNet Our Per Scene Opt.
MUSIQ($\uparrow$) 29.71 31.57 29.89 32.01
6. Conclusion